nolegirl24:

Out of the 180 students at a summer camp, 72 signed up for canoeing. There were 23 students who signed up for trekking, and 13 of those students also signed up for canoeing. Use a two-way table to organize the information and answer the following question: Approximately what percentage of students signed up for neither canoeing nor trekking? 72% 40% 54% 98% Last one and then I am done!!! ^~^

nolegirl24:

@dude This is my last one!! ^~^

dude:

We want to find the number of people who went canoeing or trekking so we add them up and subtract the 13 72+23-13= people who signed up for canoeing or trekking (an activity)

nolegirl24:

82

dude:

Right so be subtract that from the total amount because we want to find people who didn't sign up for anything

nolegirl24:

92

dude:

Right so thats the number of students, not a percentage To find percentage divide it by the total amount \(\%=\large \frac{part}{total}\)

nolegirl24:

0.51

dude:

Right so that is decimal form, to get as a percentage multiply that by 100

nolegirl24:

51

dude:

Oh wait, you subtracted incorrectly, 180-82 is not 92

nolegirl24:

o.o

nolegirl24:

its 98 my bad

dude:

Yeh

dude:

98/180

nolegirl24:

0.54 0.54*100 = 54%

dude:

Right-o

nolegirl24:

I have two more and then I am done, i missed them by accident

dude:

Its okay